Turkish finite verbs

David Pierce

2004.1.26

Contents

0 Introduction 1

1 Alphabet 1

2 Sounds 2

3 Writing 4

4 Words 4

5 Verbs: Stems 5

6 Verbs: bases 8

0 Introduction

As a student of Turkish, I make these notes in an effort to understand the logic of Turkish verbs. This is not the account of an expert. I gathered the information here mainly from the two books [1] and [2].

My account of the verbs will involve some reference to the other parts of speech. Being written, my account will use the Turkish alphabet and will therefore illustrate features like vowel-harmony that are reflected in spelling.

1 Alphabet

Like English, modern Turkish (since 1928) uses an alphabet derived from the 23-letter Latin alphabet

A, B, C, D, E, F, G, H, I, K, L, M, N, O, P, Q, R, S, T, V, X, Y, Z.

English gets 3 more letters by introducing:

• the variant J of I, and

• the variants U and W (double-U) of V.

The letter-forms given above are called upper case, to distinguish them from the lower case (a, b, c and so on).

The 29-letter Turkish alphabet can be derived from the English one by:

• eliminating Q, X and W;

• deriving from C, G, O, S and U the letters Ç, Ğ, Ö, Ş and Ü respectively, by adjoining cedilla, breve or umlaut;

• replacing I with two letters, I and İ (distinguished in both upper and lower case by whether a dot is present: the lower-case forms are therefore ı and i respectively).

The letter Ğ is called yumuşak ge (soft G), and it never begins a word. The new letters (new to English-writers, that is) can be understood as derived from the old by diacritical marks, but these marks become integral parts of the new letters. It is not the shape of the mark, but its presence at the top or bottom of the letter that is used to distinguish the letter from its ‘twin’. For example, one Turkish dairy styles itself Süta\begin{matrix}s \\ . \end{matrix} (they use a dot instead of a cedilla under the s), and there is a bus company called Uludag̀ (taking their name from the mountain at Bursa, the first capital of the Ottoman Empire; they use a grave accent rather than a breve over the g).

In the alphabetical order, I precedes İ, and the other new letters follow their ‘twins’.

Because of the distinction between the letters ı and i in Turkish, neither of these should be joined in a ligature to a preceding f. In English, the other word for a movie is printed film rather than film, but in Turkish, the word is printed as film, with four distinct letters. (Also, for example, fil means elephant or chess-bishop, while fındık means hazelnut.)

In Turkey, on a license-plate, I have seen a presumably-capital J written with a dot. But j is rare in Turkish anyway, being used only in foreign words.

2 Sounds

The name of a vowel is itself, and the name of a consonant (besides Ğ) is itself plus e.

The number of vowels is eight, that is, 2^{3}, since each vowel is determined by the quality it possesses from each of three pairs: thick/thin, flat/round and narrow/wide. These are literal translations of the usual Turkish terms

(kalın/ince, düz/yuvarlak, dar/geniş), but commonly in English the

terms are:

• back/front,

• unround/round,

• close/open.

The vowels can then be tabulated:

open  close
back front back front
unround a  e  ı  i
round  o  ö  u  ü

The sound of a can be spelled in English by uh; close the mouth more to get ı. The ö and ü are as in German; their sounds are found in French as well.

Further distinctions are possible, mainly because Turkish retains borrowings from Arabic and Persian. A difference from the norm might be shown by a circumflex: so kar means snow, but k^ar, sounding something like kyahr, means profit.

The other 21 letters are consonants. Besides the ‘semi-vowel’ y, the consonants might be tabulated:

b/p v/f  m
d/t j/ş c/ç n l

z/s

g/k ğ/h  r

The rows here are intended to correspond to position of the lips and the tip of the tongue; the columns, to the flow of breath used to pronounce the consonant. (This particular table is my own invention though, and is not the result of careful phonological study.) The members of the pairs ∗/∗ are voiced and unvoiced respectively, or soft and strong in Turkish (yumuşak and sert). The main effect of ğ is to lengthen the preceding vowel. (So it functions like gh in English in changing fit into fight.) The letter j is pronounced as in French. The sound of c is spelled in English by j or dge; the sounds of ç and ş, by ch and sh.

3 Writing

As in English, written Turkish is divided into sentences, and these into words. Some words are enclitic, being unaccented, but causing the previous syllable to be accented. (Syllabic accents are not very strong in Turkish though.) The vowel in an enclitic may also change according to the preceding vowel. This same vowel harmony is shown by many suffixes.

I shall indicate the variability of vowels with the following symbols:

• @ for an open unround vowel (a or e);

• # for a close vowel (ı, i, u or ü).

I have not seen such symbols used elsewhere in this way. (Perhaps using æ instead of @ would be more logical, but its appearance seems more confusing.) Each symbol resolves to a vowel agreeing as far as possible with the preceding vowel. In particular, after @, the only possibilities for # are ı and i.

The variability in consonants will not be indicated. However, when it begins a suffix appended to a word ending in an unvoiced (that is, strong) consonant, the letter d is unvoiced (‘strengthened’) to t. Also, terminal k changes to ğ when a suffix beginning with a vowel is added. (These changes affect the suffix -d#r and the termination -k mentioned below. Other such changes can occur.)

Example. The question Avrupa + l# + l@ş + d#r + @m@ + d#k + l@r
+ #m#z + d@n m# + s#n#z? resolves to Avrupalılaştıramadıklarımız-

dan mısınız? meaning Are you one of those whom we could not Europeanize?.

The symbol # may resolve to zero (nothing) when it begins a suffix placed after a vowel; alternatively, a consonant may be interposed, n, s or y, depending on the situation. Sometimes the y narrows the preceding @ to #.

4 Words

Not every word appears in a dictionary, even if the word is well-formed; the word may be inflected (or otherwise derived) from a dictionary-word according to standard rules. Otherwise, we might say that the same word may be used in many forms, only one or a few of which appear in the dictionary. For example, in English, we can say that man and men are two words, or else that they are two forms of the same word. Inflected forms are few in English; in Turkish they are many.

Turkish words fall into some of the same classes (parts of speech) as English words: noun, pronoun, adjective, verb and verbal noun, and adjective. Instead of prepositions, but fulfilling some of the same functions, Turkish has postpositions.

Example. The postposition gibi corresponds to the preposition like. With the noun buz ice is formed the phrase buz gibi like ice, that is, ice cold (a slogan on signs in Turkey advertising a certain American soft drink marketed all over the world).

This reversal of order in passing from the one language to the other occurs in other ways. A Turkish word (as in the earlier example) may have parts which, in English, would appear as separate words in the opposite order. As in English though, adjectives generally precede the nouns they modify.

5 Verbs: Stems

The dictionary-form of a verb is usually the infinitive, analyzed as

stem + m@k;

otherwise, the stem itself is the dictionary head-word.

All verbs are regular, except the defective verb with stem i- corresponding to the English be when used as a copula (as in ‘Socrates is a man’). The verb is defective because it does not have as many inflexions as a regular verb; it can’t have as many inflexions, because all of its inflexions are used in inflexions of regular verbs. Where there is no inflexion of i-, but the corresponding meaning is needed, then the stem ol- supplies the inflexion. Otherwise, olmak is a regular verb meaning to become. (An exception will be noted later; it is connected to the absence in Turkish of a verb meaning have.)

Every regular stem can stand by itself as a second-person singular imperative verb.

Example. The infinitives gelmek to come and çıkmak to come out or up reveal the stems gel- and çık-, hence the commands Gel! Come! and Çık! Come out! The command Be quick! is given by Çabuk ol!.

The stem of a regular verb can be understood as a simple stem together with one or more (or no) extensions. A simple verb-stem might be a verbal root, or it might be a noun or adjective with a verbal suffix, often -l@.

Example. In addition to those in the preceding example, some verbal roots include uyu- sleep and kalk- get up. Also:

• from baş head comes the stem başla- begin (make a head);

• from temiz clean comes temizle- clean (make clean);

• from köpek dog comes köpekle- cringe (make like a dog).

Extensions in a verb-stem can be seen to indicate distinctions of voice (though this means that Turkish has more voices than the active and passive of English; English does have other ways of conveying the additional distinctions of voice in Turkish).

When extensions are present, they appear in the following order:

1. reflexive: -#n;

2. reciprocal: -#ş;

3. causative:

• -t after polysyllabic stems in -l, -r or a vowel, and

• -d#r in other cases, except:

• -#r, -@r or -#t occur after some monosyllabic stems, and

• some exceptional forms occur as well;

4. passive: -#l, after stems ending in a consonant other than -l (otherwise it has the same form as the reflexive).

The significance of an extension is not always obvious from its name. Like the ancient-Greek middle voice, the reflexive extension may suggest

a doing for and not just to oneself, and it may have the same effect (and even the same form) as a passive extension (although these two kinds of extensions can be used together as well).

Used together, the reciprocal and causative extensions make the repetitive extension -#şt#r (other suffixes are used to the same effect in a few cases).

Sometimes the same verbal root can be used transitively or intransitively: gez- can mean either tour and inspect, or walk. Also, the causative extension can make an intransitive stem transitive: öl- means die, while öldür- means kill, and kayna- and kaynat- mean boil: respectively, what water does in a pot on the stove, and what I do when I put the pot there.

Used with a transitive stem, the causative extension could be called factitive: öldürt- means have [somebody] killed. As here, more than one causative extension can be used.

Here are some more examples of distinctions of voice:

Example. We have bulaşık yıka- wash dishes, but yıkan- wash oneself or be washed; the latter meaning is also expressed by yıkanıl-. Also:

• from bul- find come buluş- meet and bulun- be present;

• from ara- seek comes araştır- investigate;

• uyan- awake, uyandır- awaken, uyandırıl- be awakened.

The two-fold distinction in English between affirmation and denial (between I did it and I didn’t do it) can in Turkish be seen as three-fold. Depending on the presence of the indicated suffixes, a stem can be:

• positive: —;

• negative: -m@;

• impotential: -@m@.

In any of these cases, the stem can be further rendered potential with -@bil. These suffixes all come after any extensions.

Example. The defectiveness (in particular, the lack of infinitives) of the English modal auxiliaries gives various possibilities for translation. The second column below gives English infinitives without to (as might follow the words I shall); the third column gives a finite form (as might follow I).

  gel-  come

gelme-  not come  do not come
geleme-  not be able to come  cannot come
gelebil-  be able to come  can come
gelmiyebil-  be able not to come  may not come
gelemiyebil- possibly be unable to come may be unable to come

The potential form can be considered as a compound with the verb bil-, meaning know. Other compounds, less common, are possible; for example, compounds with dur- stop or gel- indicate continuous action. Such compounding might be said to indicate distinctions of aspect.

Moreover, simple negation is achieved, as noted, with -m@. The impotential suffix is the negation of the obsolete verb u- be powerful or able.

The Turkish verb does not make the distinction between permission and ability that English can make with may and can.

6 Verbs: bases

To a stem is added one (or sometimes more) of several characteristics (including the ‘empty’ characteristic, used, as noted, to form imperatives). The result is a base. To a base can be added personal verb-endings, or more complicated forms.

The characteristics can be classified according to:

• which of the four series of personal endings they can take, and

• whether they can be attached to the stem i-.

Their names will suggest their use in indicating distinctions of tense, aspect and mood.

Like English, Turkish has three persons and two numbers. In Turkish, these features are displayed by pronouns and verbs—also by nouns, in the sense that nouns can be given possessive suffixes that display these features.

Listed in the order 1st, 2d, 3d person singular; 1st, 2d, 3d person plural, the personal pronouns, possessive suffixes, and verb-endings of the four types are as follows.

   pronoun  ben    sen      o    biz        siz     onlar
   suffix  -#m    -#n  -(s)i  -#m#z      -#n#z     -l@r#
   type I  -#m   -s#n (-d#r)  -#z     -s#n#z  -(d#r)l@r
   type II  -m     -n  —  -k       -n#z      -l@r
   type III -@y#m  -@s#n     -@  -@l#m    -@s#n#z     -@l@r
   type IV  — -s#n         -#n, -#n#z   -s#nl@r

The type-I suffixes are said to come from the possessive suffixes (except for the parenthetical d#r, discussed below). The possessive suffixes drop the initial vowel after a vowel, and in the third-person singular, the s is dropped after a consonant.

After the possessive suffix if it is present, a noun takes a case-ending, namely one of the following:

• absolute: —,

• definite-accusative: -(y)#;

• genitive: -(n)#n;

• dative: -(y)@; • locative: -d@;

• ablative: -d@n.

The parenthetical consonants are used after vowels. The absolute case is used for subjects and indefinite direct objects.

Example. From the (originally Arabic) kitap book, and from oku- read, we can form Kitap oku Read a book, but Kitabı oku Read the book.

Any noun indicating something that is possessed needs a possessive suffix, even if the possessor is also expressed by another noun.

Example. Since mother is anne and father is baba, my mother’s father is

annemin babası.

Since there is no Turkish verb for have, the fact of possession is indicated by the (uninflected) word var existent, and in some instances by ol-.

To a noun (or pronoun or adjective) with a case-ending, a type-I ending can be added as a copula.

Example. The second-person plural here can be used (as in French) for a polite singular:

ev  house
eviniz  your house
evinizde  at your house
Evinizdeyim. I am at your house.

As proper English despises a double negative, so Turkish (at least sometimes) avoids a double plural. The ending -l@r is appended (before a possessive suffix or case-ending) to nouns, to adjectives acting as nouns, and to third-person verbs to indicate a plural; but if plurality is indicated by a number, say, than the noun remains singular in form; and a third-person verb with a plural subject might remain singular, especially if the subject is inanimate.

Example. başlar heads, beş five, so beş baş five heads (or head).

Parenthetical in the table of pronouns and endings, the suffix -d#r, unrelated to the causative extension, is said to come, as early as the 11th century, from the aorist verb turur stands. Its use is imprecisely summarized this way: in can be left out if ambiguity will not result, and it can be used for emphasis (even in persons besides the third).

Example. Türkler or Türktürler They are Turkish, but Türklerdir They are the Turks.

The characteristics are:

• (used with type-I endings)

– present: -#yor (from the ancient aorist yorır go or walk);

– future: -@c@k;

– aorist (not used with negative stems):

∗ positive: -@r, -#r or -r;

∗ negative: -m@z;

– inferential past: -m#ş;

– necessitative: -m@l# (the verbal-noun suffix -m@ with the adjectival ending -l#);

• (used with type-II endings)

– simple past: -d#;

– conditional: -s@;

• (used with type-III endings)

– subjunctive: -@ (this is given as the characteristic so that the subjunctive base will be identical with the third-person singular verb; but this vowel is just the vowel that begins each of the type-III endings);

• (used with type-IV endings)

– imperative: —.

As noted, strictly the subjunctive has no characteristic. Let us now be strict. Then the presence of a characteristic can be said to distinguish the indicative moods from the non-indicative. (By this account then, a conditional verb is indicative; this makes sense if, with the symbolic-logicians, one takes If A then B to mean just B or not-A. That the subjunctive and the imperative should be considered as allied in Turkish is suggested also by the claim that the first-plural subjunctive (type-III) ending is historically an imperative ending.)

A regular verb-base is a combination

regular verb-stem + (characteristic).

A complete finite verb can be obtained from such a base by affixing an appropriate personal ending. In particular, the indicative bases can stand alone as third-singular verbs (as the non-indicative base can stand alone as a second-singular imperative).

Example. On oku- read are formed the following indicative bases:

okuyor-  (is) reading

             okuyacak-  will read

okur-  read(s)
okumaz-  do(es) not read
okumuş-  [they say that]… w(as) reading

             okumalı-  must read

okudu-  [I saw that]… did read
okusa-  if … read(s)

In the absence of a verb, the type-I personal endings can, as noted, be suffixed to nouns, pronouns, adjectives—also to the word değil, which negates a sentence that lacks a regular verb.

Example. From uzun long, tall come Uzunum I am tall and Uzun değilsin You are not tall.

The bases formed on the stem i- are simple or compound. The simple bases result from suffixing a characteristic -m#ş, -d# or -s@ only. The nonconditional simple bases can be rendered conditional by being compounded with that simple base. Thus the bases in i- are five:

                        imiş-  imişse
                        idi-   idiyse-

                        ise-

All verb-bases have now been described. The bases formed on i- must be completed with personal endings. Personal endings may be attached to regular bases, as noted. Alternatively, a complete verb formed on a simple base in i- can follow or be suffixed to a regular verb-base.

Example. By appending bases in i- to indicative stems in oku-, these forms can be produced:

okuyormuş    okuyordu    okuyorsa    okuyormuşsa    okuyorduysa
okuyacakmış  okuyacaktı  okuyacaksa  okuyacakmışsa  okuyacaktıysa
okurmuş      okurdu      okursa      okurmuşsa      okurduysa
okumazmış    okumazdı    okumazsa    okumazmışsa    okumazdıysa
okumuş imiş  okumuştu    okumuşsa    okumuş imişsa  okumuş idiyse
okumalıymış  okumalıydı
             okuduydu    okuduysa                   okudu idiyse
okusaymış    okusaydı

If it is not associated in this way with a regular verb-base, then a complete verb in i- must follow or be suffixed to a noun, pronoun, adjective or değil.

In a verb with two characteristics, the second being conditional, the personal ending may precede this.

The distinction between the present and the aorist characteristics is of aspect. The aorist indicates habitual action; the present, an action ongoing or envisaged. A third way to indicate a present tense is by appending type-I endings or a verb in i- to an infinitive in the locative case: the result indicates ongoing action only.

Example. From yaz- write we can form:

   • Romanlar yazarım, I write novels;

• Roman yazıyorum, I am writing a novel [now, or in the future];

• Roman yazmaktayım, I am engaged in writing a novel.

The inferential characteristic is originally a past-participial ending. Used in a regular base, it indicates past action. This action is inferred from present appearances, or reported by somebody else—unless the regular base is followed by -d#r or a verb in i-, in which case no suggestion of hearsay or inference is meant. In particular, the ending -m#şt#r has the same force as -d#. The form imiş- may not indicate past time at all, but only hearsay.

Example. Barış means peace; so:

• A newspaper headline could read Barış geldi;

• The story under that head might say Barış gelmiştir;

• Somebody who read the story might report Barış gelmiş, or possibly Barış gelmiş imiş, Peace is said to have come.

The pluperfect tense is usually formed with -m#ş (+) idi-, more rarely

with -d# (+) idi.

The interrogative particle is m#; it always starts a new word. It makes a sentence into a question with a evet/hayır (yes/no) answer, and it follows the word that could also be an answer to the question. If this word is the verb, then the base is what is followed by m#; the personal ending, or the form of i- in use, is attached to this.

Exceptions include the following. If the personal ending is -l@r, then m# follows this, unless -d#r is also used. The characteristic d# of the simple past is inseparable from an attached personal ending, so if there is one, then m# must follow this.

References

[1] G. L. Lewis, Turkish Grammar, Oxford, 1967.

[2] Mehmet Hengirmen, Turkish Grammar For Foreign Students, Ankara: Engin, 2000.